# Apples to Apples: Establishing Comparability in Knowledge Generation Tasks Involving Users

### v1

@@ Abstract
Knowledge graph construction (KGC) from (semi-)structured data is challenging, and facilitating user involvement is an issue frequently brought up within this community. We cannot deny the progress we have made with respect to (declarative) knowledge generation languages and tools to help build such mappings. However, it is surprising that no two studies report on similar protocols. This heterogeneity does not allow for a comparison of KGC languages, techniques, and tools. This paper first analyses the various studies that report on studies involving users to identify the points of comparison. These gaps include a lack of systematic consistency in task design, participant selection, and evaluation metrics. Moreover, there needs to be a systematic way of analyzing the data and reporting the findings, which is also lacking. We thus propose and introduce a user protocol for KGC designed to address this challenge. Where possible, we draw and take elements from the literature we deem fit for such a protocol. The protocol, as such, allows for the comparison of languages and techniques for the RDF Mapping Languages core functionality, which is covered by most of the other state-of-the-art techniques and tools. We also propose how the protocol can be amended to compare extensions (of RML). This protocol provides an important step towards a more comparable evaluation of KGC user studies.
Resource type:
Evaluation method, protocol
License:
CC BY-SA 4.0
DOI:
URL:
@@ 1 Introduction
We preach to the choir that knowledge graphs are essential for meaningfully organizing and representing information in various domains. However, as knowledge graphs grow in complexity, efficient methods for their generation are crucial. When dealing with the challenges of (semi-)structured data sources, such as the lack of explicit semantics, which need to be aligned with ontologies or vocabularies, creating such mappings becomes a knowledge engineering task where user involvement is crucial. Users bring the necessary domain expertise to ensure the mappings are appropriate.
@@
This will enable researchers to compare different knowledge graph construction (KGC) languages and techniques, leading to a better understanding of their strengths and weaknesses and, ultimately, to more effective tools for KGC.
The contributions of this paper are twofold. First, we have a review of user studies in the KGC domain, which clearly indicates the aforementioned challenges. And the protocol we present in this paper.
Section 2 provides an overview of the related work on (declarative) approaches to KGC by mapping data sources onto RDF datasets, focusing on those that explicitly report on user studies. The goal is to show that no two papers adopt the same protocol, which makes comparing studies impossible.
Section 3 presents the protocol, which we have made available with CC-BY-SA 4.0 license. The protocol provides detailed guidelines for recruiting participants and disclosing potential biases. The process guidelines for informed consent, pre-questionnaires, familiarization activities, task execution, and post-questionnaires. The tasks consist of five sub-tasks of which, when comparing two groups, the last two can be changed to ensure a common base for comparison.
The related work will show that reporting is often limited to simple metrics and averages. Still, we deem it important to analyze the relationships between task execution, perceived usefulness, and perceived cognitive load. To this end, Section 4 proposes the statistical means to use when adopting this protocol.
In Section 5, we discuss the resources from various aspects, such as the scientific and technical, to elaborate on the soundness of our approach. This section also discusses some of the limitations. Section then concludes the paper and proposes some future directions.
@@ 2 Related Work
In [3], the authors presented an excellent survey on declarative KGC tools to help the community and practitioners choose which languages, tools, or techniques fit their needs. However, the article looks at those from a technical perspective. They look at the functionalities offered by the different options. In [2], the authors proposed a benchmark to compare KGC tools and applied it to some well-known implementations such as RMLMapper [11], Morph-KGC [1], and SDM-RDFizer [20]. It is surprising to see that the perceptions of users and practitioners have yet to be examined in a systematic manner.
@@
From a broader perspective, [23] describes three "personas" that engage with KGs: KG builders, KG analysts, and KG consumers, which were distilled from interviews with practitioners. As the name intuitively implies, the KG builder persona would be responsible for generating the KG from heterogeneous sources, but the persona is also in charge of ontology engineering. The authors state that builders could benefit from tools that help them ensure that the schema is respected (what the authors call an "enforcer") as well as adequate visualization tools. While the paper does not explicitly mention KG generation and mappings as tasks, they fall under the "data integration" umbrella. Key is here is that the interviews indicate that there are challenges impeding uptake.
@@
It seems that the role of practitioners, or users, is at times neglected. This is certainly the case for KG Generation, as we will now demonstrate via our literature review. Our review looked at the following papers reporting on users, their experiences, and/or perceived usability: [25, 17, 26, 4, 9, 16, 8, 7, 13, 28, 6]. Most of these studies looked at the creation of mappings. Exceptions are [7] reporting on studies on mapping understanding and [6] reporting on a complex data flow that included mapping creation. [^1] Excluded from this survey are publications that do not report on users. For example, in [18], the authors reported involving participants, but no report on the participant’s experience was made. Other examples of papers mentioning users, participants, etc. without any detailed reporting include [19], [15], and [21]. We compare the various aspects of these user studies in Tables 1, 2, and 3. From these tables, we can observe a couple of important points:
  • Some report on comparing mapping languages and/and tools (e.g., [16] and [8]), and others report on comparing mapping languages (e.g., [13] and [28]). Quite a few papers merely report on the perceived usability of their tool without any comparison. We argue that reporting on user studies only makes sense if there is a basis for comparison.
@@
  • Looking at the procedure, we see many recurring elements (some (training) resources are being shared, pre-assessment surveys, introduction of tasks, surveys, etc.). No two procedures are the same, which hampers our ability to compare the studies. Some studies reported asking about information such as gender and age but did not report on those in the data analysis.
@@
From this survey, we can conclude that there is a critical need for homogeneous protocols, including tasks, for comparing advances in KGC approaches (mapping languages and tools alike). In the next section, we propose a protocol to address these issues and how they can be used.
Table 1: Comparison of KGC languages, techniques, and tools in which participants were observed. We can observe that procedures vary widely.
Table 2: Comparison of KGC languages, techniques, and tools in which participants were observed. We can observe that no two papers share tasks and data analysis techniques.
Table 3: Comparison of KGC languages, techniques, and tools in which participants were observed. We can observe self-reported prior knowledge and skills and participant selection varies widely. Also, the number of participants varies.
3 The KGC User Study Protocol
@@
The protocol as such can be used to analyze the perceived usability and cognitive load of a mapping language, tool, … as well as the accuracy achieved by users and the task execution time. When users use this protocol on only one group, the hypotheses are limited to comparing participants with different demographics, or by comparing the results with other experiments using this protocol. Scholars adopting this protocol can easily extend the protocol to compare two tools, techniques, or languages. This will be explained in Section 3.5.
As we will explain in Section 3.5, there tasks can
@@ 3.1 Participant Selection
Adopters of the protocol should indicate how participants are recruited and from where. Adopters should disclose potential biases by providing details about factors influencing the study or participant behavior. Examples include the hierarchical relationship between research group leaders and researchers and students recruited from classes. There is also a difference between voluntary participation and mandatory representation (e.g., in the context of a teaching activity).
There is an unwritten rule that states that 5 participants are sufficient to get an idea of a system’s perceived usability or usefulness, but, as [12] observed in an experiment, "increasing the number from 5 to 10 can result in a dramatic improvement in data confidence." They also found that increasing the number to 20 practically guarantees all problems to be seen, but we recognize that recruiting participants is difficult. As such, we require 10 participants per group.
@@ 3.2 Process
Participants begin by reviewing informed consent materials and completing a pre-questionnaire assessing their demographics, prior knowledge, and expectations. Next, they attend a presentation introducing the technology, review relevant documentation, and engage in a familiarization activity. The core of the study involves participants executing a defined task using the technology. Finally, participants complete a post-questionnaire evaluating their experience, including usability, efficiency, and perceived cognitive load. This structured process aims to gather comprehensive data on user interaction and perception of the technology.
@@
Furthermore, we ask authors to report on how responses were submitted (e.g., email, paper, form) and the anticipated duration of the experiment. Note that while in-class experiments often have time constraints, other environments may be more flexible. Our protocol foresees 1 hour for the five tasks. If, for example, all steps are conducted in a classroom setting, the experiment would require 2:30. Finally, clarify whether participants will be allowed to ask questions during the experiment and help should be limited to aspects not core to the KG generation process and experiment. For example, helping participants navigate to the correct directory in a terminal or assessing whether there is a network issue is permitted, but providing help on how to execute a mapping is not. Ideally, studies would report on those (and their number of occurrences).
@@ 3.3 Pre-questionnaire
Studies often inquire about participants and group them based on self-reported information on their background and proficiency with specific techniques. We aim to homogenize this by proposing an exhaustive list of current roles, formal training, and self-perceived competency levels in certain Semantic Web technologies. We also included three questions related to intrinsic motivation (enjoyment, curiosity, and value). It is important to analyze the impact of self-selection bias in voluntary participation.
@@ 3.4 Mapping tasks
Participants will be requested to complete five mapping tasks, some of which are interdependent. In our state-of-the-art surveys, many different types (e.g., mapping understanding vs. mapping creation) are applied in many domains. We propose choosing a general domain for participants to understand: projects, project tasks, and employees who manage projects and are assigned tasks. The rationale is that there are three distinct types, and users should, at this point, not struggle with ontological questions such as roles vs. types.
Figure 1 depicts the UoD of the data to be transformed into RDF. We use an ERD, but JSON and XML files can easily represent the data. To ensure attribute names do not mislead participants, we ensured all attributes are unambiguous. In this simple UoD, all relations are many-to-one, though this can be easily extended to many-to-many when transforming documents.
Figure 1: Enter Caption
@@
    Generate ex:of and ex:assignedTo properties from tasks to, respectively, projects and employees.
@@
We draw attention to the fact that employees’ IRIs are based on their names, whereas the project data sources refer to employees via their IDs. This requires users to join the data in the two sources. In the case of RML, this requires users to "join" the two sources either at the level of the logical source or by using a referencing object map.
@@ 3.5 Variants
We stated that the tasks focus on RML-core functionality, which is also covered by other KGC languages and techniques such as ShExML [13] and SPARQL Anything [10]. While one can argue that the set of desired functionalities is limited, [21] covered the requirements of all RML extensions. It is not feasible to formulate tasks that cover practically all use cases, not only in time but also in data source complexity.
In this section, we describe how this protocol can be used for comparing different aspects.
  • When comparing languages, techniques, or tools. One can provide the same tasks to two different groups. One can compare different mapping languages (e.g., ShExML vs. RML), compare editors vs. "bare bones" languages (e.g., RMLEditor vs. RML), compare languages and abstractions of languages (e.g., RML vs. YARRRML), and even different editors and languages.
  • When the aim is to compare support for an advanced KGC requirement such as named graphs, collections and containers, or RDF* (among others), then one can take this protocol as is for one group, and only change the last two tasks for the second in which those requirements are covered, with the first three tasks giving a comparable baseline, unless the extension covered another aspect such as storing the resulting trip.
  • When the aim is to such compare KG construction across languages, techniques, or tools, then the protocol must be amended for both. Again, the first three tasks must remain unchanged as to ensure some comparative baseline with literature. This approach can be used to compare the languages and tools for more advanced KGC requirements such as named graphs, RDF collections and containers, and function calls, among others.
Participants are assumed to have access to prepared "resources" or "environments" to focus on the tasks. In the case of RML, for instance, the logical sources would be provided in the tool or for them to copy and paste. This allows researchers to assess the languages and tools with respect to these aspects by giving one group the prepared artifacts and requesting the other to formulate the logical sources themselves. We deem this a special case of comparing a baseline with an extension as described above, but where the five tasks remain unchanged.
@@ 3.6 Post-questionnaire
Both SUS and PSSUQ are used to measure usability. SUS is adequate for a rapid and general measure of a system’s usability. Still, the latter offers more advantages because it assesses three aspects of a system: system usefulness, information quality, and interface quality. Furthermore, there is a question about the system as a whole, which allows one to damped the perception of individual aspects. The original PSSUQ survey uses 19 questions (as adopted by [9], for instance), but recent iterations have removed three redundant questions.
@@
  • NASA-TLX has been validated in several domains [14] and combines six factors believed to influence the mental workload. Each factor is quantified with a subjective judgment and a weight computed via a paired comparison procedure. For each possible pair of the six factors, participants must decide which factor contributed the most to the mental workload during the task. The weights $w$ are the number of times each dimension was selected. The possible weights range from 0 (irrelevant) to 5 (most important). The final score is computed as a weighted average, considering the subjective rating of each attribute $d_{i}$ and the correspondent weights $w_{i}$: $\frac{1}{15}\sum_{i=1}^{6}{d_{i}*w_{i}}$. It is possible to calculate the scores by eliminating the weighted procedure, which yields the so-called Raw TLX.
@@
  • Accuracy should be determined by graph isomorphism (did they generate the expected graph, which is true or false), and, for more nuanced numbers, precision (the proportion of triples that are generated that are in the expected graph), recall (the proportion of expected triples that are in the generated graph), and F-measure combining precision and recall.
This approach accounts for situations where a participant generates additional triples, for instance. Figures should be reported both on task and global level.
  • Efficiency is often erroneously conflated with task execution time. The former is a broad concept that measures how well resources are used, which is not only time. Most studies measured the time it took for tasks to be completed. We require studies to report on task execution time, and the method to measure it. One can manually record time or record user interactions to time the tasks retroactively. Another approach is to request users to report on the time or use electronic forms that keep track of time.
@@
As the experiment should not be too time-consuming, we avoided interviews to obtain qualitative feedback. We also avoid "think aloud" experiments as they can impact the cognitive load. Whether a study reports on it or not, we require studies to report on any additional instruments they used. There is, however, a qualitative dimension to our protocol, as the PSSUQ does leave room for comments on each of the 16 questions.

### v2

# A Protocol for KG Construction Tasks Involving Users

@@ Abstract
Knowledge graph construction (KGC) from (semi-)structured data is challenging, and facilitating user involvement is an issue frequently brought up within this community. We cannot deny the progress we have made with respect to (declarative) knowledge graph construction languages and tools to help build such mappings. However, it is surprising that no two studies report on similar protocols. This heterogeneity does not allow for comparing KGC languages, techniques, and tools. This paper first analyses studies involving users to identify the points of comparison. These gaps include a lack of systematic consistency in task design, participant selection, and evaluation metrics. Moreover, there needs to be a systematic way of analyzing the data and reporting the findings, which is also lacking. We thus propose and introduce a user protocol for KGC designed to address this challenge. Where possible, we draw and take elements from the literature we deem fit for such a protocol. The protocol, as such, allows for the comparison of languages and techniques for the RDF Mapping Language (RML) core functionality, which is covered by most of the other state-of-the-art techniques and tools. We also propose how the protocol can be amended to compare extensions (of RML). This protocol provides an important step towards a more comparable evaluation of KGC user studies.
@@ 1 Introduction
We preach to the choir that knowledge graphs are essential for meaningfully organizing and representing information in various domains. However, as knowledge graphs grow in complexity, efficient methods for their construction (or generation) are crucial. When dealing with the challenges of (semi-)structured data sources, such as the lack of explicit semantics, which need to be aligned with ontologies or vocabularies, creating such mappings becomes a knowledge engineering task where user involvement is crucial. Users bring the necessary domain expertise to ensure the mappings are appropriate.
@@
This will enable researchers to compare different knowledge graph construction languages and techniques, leading to a better understanding of their strengths and weaknesses and, ultimately, to more effective tools for KGC.
This paper’s contributions are twofold. First, we review user studies in the KGC domain, which indicate the abovementioned challenges. The second contribution is the protocol. Section 2 provides an overview of the related work on (declarative) approaches to KGC by mapping data sources onto RDF datasets, focusing on those that explicitly report on user studies. The goal is to show that no two papers adopt the same protocol, which makes comparing studies impossible. Section 3 presents the protocol we have made available with CC-BY-SA 4.0 license. The protocol provides detailed guidelines for recruiting participants and disclosing potential biases. The process guidelines for informed consent, pre-questionnaires, familiarization activities, task execution, and post-questionnaires. The tasks consist of five tasks, of which, when comparing two groups, the last two can be changed to ensure a common base for comparison. The related work will show that reporting is often limited to simple metrics and averages. Still, we deem it important to analyze the relationships between task execution, perceived usefulness, and perceived cognitive load. To this end, Section 4 proposes the statistical means to use when adopting this protocol. In Section 5, we discuss the resources from various aspects, such as the scientific and technical, to elaborate on the soundness of our approach. This section also discusses some of the limitations. Section 6 then concludes the paper and proposes future directions.
@@ 2 Related Work
In [1], the authors presented an excellent survey on declarative KGC tools to help the community and practitioners choose which languages, tools, or techniques fit their needs. However, the article looks at those from a technical perspective. They look at the functionalities offered by the different options. In [2], the authors proposed a benchmark to compare KGC tools and applied it to some well-known implementations such as RMLMapper [^1] https://github.com/RMLio/rmlmapper-java, Morph-KGC [3], and SDM-RDFizer [4]. It is surprising to see that the perceptions of users and practitioners have yet to be examined in a systematic manner.
@@
From a broader perspective, [5] describes three “personas” that engage with KGs: KG builders, KG analysts, and KG consumers, which were distilled from interviews with practitioners. As the name intuitively implies, the KG builder persona would be responsible for generating the KG from heterogeneous sources, but the persona is also in charge of ontology engineering. The authors state that builders could benefit from tools that help them ensure that the schema is respected (what the authors call an “enforcer”) as well as adequate visualization tools. While the paper does not explicitly mention KG construction and mappings as tasks, they fall under the “data integration” umbrella. The interviews indicate that there are challenges impeding uptake.
@@
It seems that practitioners’ or users’ roles are sometimes neglected. This is certainly the case for KG construction, as we will demonstrate via our literature review. Our review looked at the following papers reporting on users, their experiences, and/or perceived usability: [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Most of these studies looked at the creation of mappings. Exceptions are [13] reporting on studies on mapping understanding and [16] reporting on a complex data flow that included mapping creation. [^2] Excluded from this survey are publications that do not report on users. For example, in [17], the authors reported involving participants, but no report on the participants’ experience was made. Other examples of papers mentioning users, participants, etc. without any detailed reporting include [18], [19], and [20]. We compare the various aspects of these user studies in Tables 1, 2, and 3. From these tables, we can observe a couple of important points:
  • Some report on comparing mapping languages and/or tools (e.g., [11] and [12]), and others report on comparing mapping languages (e.g., [14] and [15]). Quite a few papers merely report on the perceived usability of their tool without any comparison. We argue that reporting on user studies only makes sense if there is a basis for comparison. Without a reference point (e.g., a comparable tool evaluated under a shared protocol) or even a common protocol to establish such points, it becomes difficult to interpret the significance of usability results. For instance, reporting that users found a tool “easy to use” or completed a task within a certain time frame lacks context unless these outcomes can be meaningfully compared. Comparative user studies are therefore essential to provide this community with insights and to guide the development of more effective knowledge graph construction languages and tools.
@@
  • Looking at the procedure, we see many recurring elements (some (training) resources are being shared, pre-assessment surveys, introduction of tasks, surveys, etc.). No two procedures are the same, which limits our ability to compare the studies. Some studies reported asking about information such as gender and age but did not report on those in the data analysis.
@@
From this survey, we can conclude that there is a critical need for homogeneous protocols, including tasks, for comparing advances in KG construction (KGC) approaches (mapping languages and tools alike). In the next section, we propose a protocol to address these issues and how they can be used.
Table 1: Comparison of KG construction languages, techniques, and tools in which participants were observed. We can observe that procedures vary widely.
Table 2: Comparison of KG construction languages, techniques, and tools in which participants were observed. We can observe that no two papers share tasks and data analysis techniques.
Table 3: Comparison of KG construction languages, techniques, and tools in which participants were observed. We can observe that self-reported prior knowledge, self-reported expertise, and participant selection vary widely. Also, the number of participants varies.
3 The KG Construction User Study Protocol
@@
The protocol can be used to analyze the perceived usability and cognitive load of a mapping language or tool, as well as the accuracy achieved by users and the task execution time. When scholars use this protocol on only one group, the hypotheses are limited to comparing participants with different backgrounds or comparing the results with other experiments using that same protocol. Scholars adopting this protocol can easily extend the protocol to compare two tools, techniques, or languages. This will be explained in Section 3.5.
The focus of the protocol is to facilitate the discovery of problems (i.e., formative testing), not on task-level measurements. [23] describes the difference between the two. The protocol proposes task-level measurements and techniques for analyzing the data. When sample sizes are small, they can merely give insights.
@@ 3.1 Participant Selection
Adopters of the protocol should indicate how participants are recruited and from where. Adopters should disclose potential biases by providing details about factors influencing the study or participant behavior. Examples include the hierarchical relationship between research group leaders and researchers, as well as students recruited from classes. There is also a difference between voluntary participation and mandatory representation (e.g., in the context of a teaching activity), which may lead to non-probability samples. The absence of certain participants may lead to response bias, which is the possible impact on observations had those participants taken part in the experiment. In [24], the authors describe how disclosing participant selection is important to recognize potential biases.
Practitioners involved with UI and UX often state that five users are sufficient to discover most of the usability problems. That is more likely the case for problem discovery than task-level measurements, which require larger sample sizes [23]. As [25] observed in an experiment, “increasing the number from 5 to 10 can result in a dramatic improvement in data confidence.” They also found that increasing the number to 20 practically guarantees all problems to be seen, but we recognize that recruiting participants is difficult. As such, we (strongly) recommend 10 participants per group.
@@ 3.2 Process
Participants begin by reviewing informed consent materials and completing a pre-questionnaire assessing their backgrounds, prior knowledge, and expectations. Next, they attend a presentation introducing the technology, review relevant documentation, and engage in a familiarization activity. The core of the study involves participants executing a defined task using the technology. Finally, participants complete a post-questionnaire evaluating their experience, including usability, efficiency, and perceived cognitive load. This structured process aims to gather comprehensive data on user interaction and perception of the technology.
@@
Furthermore, we ask authors to report on how responses were submitted (e.g., email, paper, form) and the anticipated duration of the experiment. While in-class experiments often have time constraints, other environments may be more flexible. Our protocol foresees 1 hour for the five tasks. If, for example, all steps are conducted in a classroom setting, the experiment would require 2:30 in total (i.e., including the questionnaires and consulting the training material). Finally, clarify whether participants can ask questions during the experiment. Help should be limited to aspects not core to the KG construction process and experiment. For example, helping participants navigate to the correct directory in a terminal or assessing whether a network issue is permitted, but providing help to execute a mapping is not. Ideally, studies should report on those (and their number of occurrences).
@@ 3.3 Pre-questionnaire
Studies often inquire about participants and group them based on self-reported information on their background and proficiency with specific techniques. We aim to homogenize this by proposing an exhaustive list of current roles, formal training, and self-perceived competency levels in certain Semantic Web technologies. We also included three questions related to intrinsic motivation (enjoyment, curiosity, and value). It is important to analyze the impact of self-selection bias in voluntary participation. This allows us to explore potential self-selection bias in voluntary participation by examining correlations between motivation levels, task performance, and perceived usability. In contrast to self-selection bias in voluntary participation, mandatory settings introduce the risk of low engagement among less intrinsically motivated participants. Including questions on intrinsic motivation enables us to assess how motivation levels influence task performance and perceived usability.
@@ 3.4 Mapping tasks
Participants will be requested to complete five mapping tasks, some of which are interdependent. We can observe that different studies adopt different domains, some of which are specific (e.g., health care). We argue that participants should question the domain used in the experiment. Therefore, we propose a domain that is sufficiently generic and accessible for participants to understand. Our protocol proposes mapping data about projects, project tasks, and employees who manage projects and are assigned tasks. Figure 1 depicts the Universe of Discourse (UoD) of the data to be transformed into RDF. We use an Entity-Relationship Diagram (ERD), but JSON and XML files can easily represent the data. [^4] Examples are provided in the GitHub repository under the “assets” directory. To ensure attribute names do not confuse participants, we ensured all attributes are unambiguous. In this simple UoD, all relations are many-to-one, though this can be easily extended to many-to-many when transforming documents.
Figure 1: Entity-Relationship Diagram (ERD) representing the UoD of the data used in the protocol.
@@
    Generate ex:of and ex:assignedTo properties from tasks to, resp., projects and employees.
@@
We draw attention to the fact that employees’ IRIs are based on their names, whereas the project data sources refer to employees via their IDs. This requires users to join the data in the two sources. In the case of RML, this requires users to “join” the two sources either at the level of the logical source or by using a referencing object map.
@@ 3.5 Variants
We stated that the tasks focus on RML-core functionality, which is also covered by other KGC languages and techniques such as ShExML [14] and SPARQL Anything [26]. While one can argue that the set of desired functionalities is limited, [20] covered the requirements of all RML extensions. Formulating tasks that cover practically all possible features and scenarios, not only in time but also in data source complexity, is not feasible. However, when a particular feature or scenario needs to be investigated, the protocol can be amended. This section describes how this protocol can be adapted and used to compare different tools, features, or aspects.
  • One can provide the same tasks to two different groups when comparing languages, techniques, or tools. One can compare different mapping languages (e.g., ShExML vs. RML), compare editors vs. “bare bones” languages (e.g., RMLEditor vs. RML), compare languages and abstractions of languages (e.g., RML vs. YARRRML), and even different editors and languages.
  • When the aim is to compare a tool or language’s support for an advanced KGC requirement such as named graphs, collections and containers, or RDF-star (among others), then one can take this protocol as is for one group, and only change the last two tasks for the second in which those requirements are covered for the second group. The first three tasks, which are shared, provide a basis for comparing the two groups. This design ensures that both groups share a comparable foundation (tasks 1-3), allowing us to isolate and evaluate the impact of the new features introduced in the final tasks.
  • We have proposed a generic and accessible domain for the protocol to ensure broad applicability. However, the protocol can be adapted to include similar tasks within a different (and domain-specific) context. This adaptation would enable researchers to assess whether a language or tool designed for a particular application domain performs better in its setting. In such cases, it is important to compare the tool or language across both the original (generic) domain and the adapted (domain-specific) version.
Participants are assumed to have access to prepared “resources” or “environments” to focus on the tasks. In the case of RML, for instance, the logical sources would be provided in the tool or for them to copy and paste. This allows researchers to assess the languages and tools with respect to these aspects by giving one group the prepared artifacts and requesting the other to formulate the logical sources themselves. We deem this a special case of comparing a baseline with an extension as described above, but where the five tasks remain unchanged.
@@ 3.6 Post-questionnaire
Both SUS and PSSUQ are used to measure usability. SUS is adequate for a rapid and general measure of a system’s usability. Still, the latter offers more advantages because it assesses three aspects of a system: system usefulness, information quality, and interface quality. Furthermore, there is a question about the system as a whole, which allows one to dampen the perception of individual aspects. The original PSSUQ survey uses 19 questions (as adopted by [10], for instance), but recent iterations have removed three redundant questions.
@@
  • NASA-TLX has been validated in several domains [28] and combines six factors believed to influence the mental workload. Each factor is quantified with a subjective judgment and a weight computed via a paired comparison procedure. For each possible pair of the six factors, participants must decide which factor contributed the most to the mental workload during the task. The weights $w$ are the number of times each dimension was selected. The possible weights range from 0 (irrelevant) to 5 (most important). The final score is computed as a weighted average, considering the subjective rating of each attribute $d_{i}$ and the corresponding weights $w_{i}$: $\frac{1}{15}\sum_{i=1}^{6}{d_{i}*w_{i}}$. It is possible to calculate the scores by eliminating the weighted procedure, which yields the so-called Raw TLX.
@@
  • Accuracy should be determined by (1) graph isomorphism (did they generate the expected graph, which is true or false), and, for more nuanced numbers, (2) precision (the proportion of triples that are generated that are in the expected graph), (3) recall (the proportion of expected triples that are in the generated graph), and (4) F-measure combining precision and recall.
This approach accounts for situations where a participant generates additional triples, for instance. Figures should be reported both on the task and global levels.
  • Efficiency is often erroneously conflated with task execution time. The former is a broad concept that measures how well resources are used, which is not only time. Most studies measured the time it took for tasks to be completed. We recommend studies to report on task execution time, and the method to measure it. One can manually record time or use software to record user interactions to time the tasks. Another approach is to request users to report on the time or use electronic forms that keep track of time.
@@
As the experiment should not be too time-consuming, we avoided interviews to obtain qualitative feedback. We also avoid “think aloud” experiments as they can impact the cognitive load. Whether a study reports on it or not, we recommend studies to report on any additional instruments they used. There is, however, a qualitative dimension to our protocol, as the PSSUQ does leave room for comments on each of the 16 questions.